Use Descartes’ circle method to determine the slope of the normal line (and subsequently, the slope of the tangent line) to y = x5/2
Answer #1

Please upvote.

solution Let there is a circle with centre (9,0) and passes through the carne y apela at (6, 6512) en of circle will be cx-a²Hence (9,0) = (55*+25,0) The line from (9,0) to (6,512) is the normal and it has slope = . b 120 b-a =b- 55*725 2 Bo64-66 GIN+ + 23 4 x + x(1+8) + ( 64 6-20) 1a-b))of+x?20x-64296-55( box+x=-298-64248-65 らくスタトスト29ペーム +296-25 az1B)-298-57420665 マスキン -